# On Developers’ Self-Declaration of AI-Generated Code: An Analysis of Practices

### v1

@@ 5.2. Reasons for Self-Declaring (or Not) AI-Generated Code
Our survey results show that developers self-declare AI-generated code due to several reasons, including (1) the need for tracking and monitoring AI-generated code in code files, (2) promoting ethical use of AI-generated code, (3) providing context to share knowledge about AI-generated code and enhance collaboration, and (4) indicating the code quality by sharing whether the code is tested or undergone human modification.
Tracking and monitoring code for later reviews and debugging. A key reason developers self-declare AI-generated code is to track its usage for future code reviews, debugging, and enhancement. The self-declaration is primarily used to differentiate between AI-generated and human-written code. This distinction becomes particularly important when developers do not verify the functionality or enhance the quality of AI-generated code. There is a need for reviewing and modifying any AI-generated code, with several previous studies denoting possible quality issues in AI-generated code and suggesting continued verification and extensive testing (Jin et al., 2024; Ouyang et al., 2024).
Transparency, accountability, and ethical considerations. Another reason for self-declaring AI-generated code is that it is considered an ethical obligation. Several previous studies have explored ethics in AI, specifically in software development (Stalnaker et al., 2024; Xu et al., 2025). Nonetheless, existing research lacks clear guidelines on the ethical use of AI-generated code in practice. Developers noted that self-declaring AI-generated code facilitates transparency in the development process, and helps track its provenance (Xu et al., 2025).
Evaluating and improving code quality. Self-declaring AI-generated code allows developers to indicate whether the code is entirely AI-generated, tested, or contains human contributions. The self-declaration comments show the reliability of the code by allowing developers to assess the quality of the code and work on reviewing and improving it. This way of writing self-declaration comments aligns with previous studies, which raise concerns about the quality of AI-generated code (Liu et al., 2024a; Ouyang et al., 2024). Due to these concerns, human review and improvements are inevitable for AI-generated code to satisfy the code quality requirement (Dakhel et al., 2023).
Knowledge sharing within a team. Self-declaring AI-generated code helps developers improve knowledge sharing and collaboration by providing explanations and contextual information in the self-declaration comments. In our mining study of GitHub, we identified short code explanations and contextual information, such as prompts used to generate code. Developers include such information in their self-declaration comments to provide context for AI-generated code to other team members, sharing knowledge about the code and enhancing collaboration. Similar to our study, Yu et al. reported that developers share prompts used to generate code and sometimes provide additional details in code comments (Yu et al., 2024). Stalnaker et al. also found that developers record the prompts or the entire interaction with AI tools to document AI-generated code (Stalnaker et al., 2024).
@@
Our survey results show that there are two reasons for not self-declaring AI-generated code: (1) the extensive modifications to AI-generated code and (2) the perception of AI code generation tools as productivity enhancers rather than autonomous contributors.
Customization and modification. The main reason is that developers do not self-declare AI-generated code because the code undergoes a cycle of iterative refinement before it is added to their projects. The rationale behind this modification is twofold: (1) Developers use the generated code as a source of inspiration only. Although AI code generation tools have undergone significant improvements over the years, their output often still requires customization to meet project requirements. (2) AI-generated code may contain code quality issues that require human modification before it becomes a viable solution. As developers extensively verify and modify the code according to their requirements, the generated code evolves through additional human contributions. Therefore, developers feel that the code is now their contribution.
Self-declaration is unnecessary. Another reason is that developers do see those AI code generation tools similarly to reference materials such as online documentation and coding forums (e.g., Stack Overflow). Therefore, they view the generated code as similar to adapting examples from online forums or documentation. Developers utilize AI code generation tools to enhance productivity; however, they do not employ these tools as autonomous contributors. Therefore, they do not feel the need to self-declare AI-generated code.
@@ 5.3. Implications
Self-declaring AI-generated code, especially when the generated code is not fully understood or tested. Developers should self-declare AI-generated code, particularly when it has been generated with little human intervention, or when the generated code has not been thoroughly tested or verified. Even though the code is thoroughly reviewed, understood, and tested, self-declaration remains valuable for the ethical use of AI-generated code. In both cases, self-declaring the code is beneficial for maintaining a record of its origin and supporting future maintenance.
@@
Using simple self-declaration comments for better transparency. When adding AI-generated code to a project, developers may include a short self-declaration comment. This comment should note that the code was AI-generated, the name of the tool used, and the name of the code snippet (e.g., the name of the class, function, or language construct). This practice helps ensure transparency in development, promotes the ethical use of AI-generated code, and facilitates the distinction between AI-generated and human-written code.
@@
Adding code explanations for complex AI-generated code. A good practice for developers is to include a brief explanation alongside the simple self-declaration comments, especially when the AI-generated code is relatively complex. These comments should describe how the code works, why it is added, and any key details about the integrated AI-generated code that team members may need in the future. These code explanations should be kept concise, and developers should be mindful of the code’s readability.
@@
Include code quality indications in self-declaration comments. Developers should add details about the quality of AI-generated code in the self-declaration comments. This can include whether the code has been tested, whether it was modified by a human, and if there are any known issues or limitations in the AI-generated code. Including such information helps team members assess the code’s reliability and determine if further improvements are necessary.
@@
Document the context of AI-generated code. When using AI-generated code, developers should consider documenting its use context, such as the prompts used. Keeping a record of such information can help developers understand how the code was generated and make it easier to modify or troubleshoot in the future.
@@
Using snippet-level comments for better traceability. It is advised to use snippet-level rather than file-level comments when self-declaring AI-generated code, as they provide a precise and more detailed approach to track AI-generated code snippets. However, if a file contains mostly AI-generated code or multiple AI-generated code snippets, file-level comments may be more suitable, as they offer a more concise summary without excessive repetition.
@@ 6.2. Construct Validity
Keyword-based search. We used a set of search terms to identify and locate AI-generated code on GitHub repositories. This approach introduced a threat to validity as the search terms used might not be comprehensive or inclusive. To minimize this threat, we first conducted a pilot search using a small set of keywords, which helped us finalize the list of potential search terms. We iteratively added new terms to improve the coverage of our keyword. As a result, we identified a list of 23 search terms with a broad coverage.
@@ 7. Conclusions and Future Work
This study explores the practices of developers’ self-declaration of AI-generated code. We examined whether developers self-declare AI-generated code, their motivations for doing so, and the methods they use to achieve this. We employed a mixed-methods approach, combining data from GitHub repositories with a follow-up industrial survey among practitioners. We found that developers have varying opinions on self-declaring AI-generated code, and their self-declaration practices are influenced by the proportion of AI contribution in the generated code. Most practitioners (76.6%) always or sometimes self-declare AI-generated code. In contrast, other practitioners (23.4%) noted that they never self-declare AI-generated code. Top reasons for self-declaring AI-generated code include (1) the need to track and monitor code for future review and debugging, and (2) ethical considerations related to the use of code generation tools. On the other hand, reasons for not self-declaring AI-generated code include (1) AI-generated code often requires extensive human modification, and (2) developers view self-declaration as an unnecessary activity, comparable to utilizing online documentation or coding forums. We then proposed a set of guidelines for developers to follow when self-declaring AI-generated code.
Our study also opens up several directions for future research. (1) We identify various reasons for self-declaring (or not) AI-generated code, which can be further investigated with a larger set of developers. Additionally, we provide guidelines for self-declaring AI-generated code by analyzing how developers report their generated code based on data from the mining study and industrial survey. Future research can further refine and expand the guidelines using new datasets and validate their usefulness through empirical evaluation. (2) Our research shows that developers consider self-declaration comments beneficial for tracking AI-generated code for future review, debugging, and improvement. This opens up avenues for future research to investigate the impact of self-declaration comments on various code quality attributes, such as readability and maintainability. Such studies can contribute to a deeper understanding of the evolving landscape of software development, where AI-generated code is increasingly integrated into modern software systems. (3) It would be valuable to provide a code-commenting function in AI code generation tools to support the guidelines (see Section 5.3) for developers to declare AI-generated code.

### v2

@@ 5.2. Reasons for Self-Declaring (or Not) AI-Generated Code
Our survey shows that the primary reasons developers self-declare AI-generated code are the need for tracking and monitoring AI-generated code, as well as ethical considerations. Existing literature has extensively highlighted the quality and security risks associated with AI-generated code, emphasizing the importance of reviewing and improving it by experienced human developers (Jin et al., 2024; Ouyang et al., 2024). This highlights the need for accurately tracking AI-generated contributions to support future code review, debugging, and maintenance. While several studies have explored automated techniques for detecting AI-generated code (Gurioli et al., 2025; Li et al., 2023), these approaches face challenges in identifying AI-generated code accurately. Interestingly, developers also shared in their responses that they integrate AI-generated code after modifying and improving it. This modification can further reduce the accuracy when detecting AI-generated code using automatic approaches. This highlights the importance of self-declaring AI-generated code for accurately tracking it as suggested by survey participants. Additionally, developers considered self-declaring AI-generated code as an ethical obligation. This shows developers’ concern for ethical AI-assisted software development and responsible human-AI collaboration. Developers noted that self-declaring AI-generated code facilitates transparency in the development process and helps track its provenance. Several previous studies have explored ethics in AI, specifically in software development (Stalnaker et al., 2024; Xu et al., 2025). Nonetheless, existing research lacks clear guidelines on the ethical use of AI-generated code in practice.
@@
The main reason is that developers do not self-declare AI-generated code because the code undergoes a cycle of iterative refinement before it is added to their projects. The rationale behind this modification is twofold: (1) Developers use the generated code as a source of inspiration only. Although AI code generation tools have undergone significant improvements over the years, their output often still requires customization to meet project requirements. (2) AI-generated code may contain code quality issues that require human modification before it becomes a viable solution. As developers extensively verify and modify the code according to their requirements, the generated code evolves through additional human contributions. Therefore, developers feel that the code is now their contribution. Another reason is that developers view these AI code generation tools similarly to reference materials, such as online documentation and coding forums (e.g., Stack Overflow). Therefore, they view the generated code as similar to adapting examples from online forums or documentation. Developers utilize AI code generation tools to enhance productivity; however, they do not employ these tools as autonomous contributors. Therefore, they do not feel the need to self-declare AI-generated code.
@@ 5.3. Implications
5.3.1. Implications for Researchers
Our findings highlight several implications for the future of AI-assisted software engineering. (1) The low frequency of self-declaration comments suggests the need for automated tools that can detect AI-generated code and insert structured self-declaration comments with the developer’s confirmation to reduce manual effort while improving traceability. Such tools could also allow the attachment of useful additional information, such as prompts and quality indications, to support debugging and code review. (2) The use of snippet-level and file-level comments observed in our study indicates that future IDEs should provide flexible mechanisms for comments to self-declare AI-generated contributions at different levels of granularity. (3) The inclusion of quality indicators in self-declaration comments shows potential for integrating automated quality checks with self-declaration comments (e.g., attaching a negative indicator to a self-declaration comment if unit tests or static analysis fail). This feature will enable teams to better assess the reliability of AI-generated code. (4) Developers’ ethical motivations for self-declaration indicate the importance of establishing organizational guidelines and industry standards for AI-generated code provenance, while also creating opportunities for future research to explore how automated self-declaration can be integrated into existing development workflows.
5.3.2. Implications for Practitioners
@@
Self-declaring AI-generated code, especially when the generated code is not fully understood or tested. Developers should self-declare AI-generated code, particularly when it has been generated with little human intervention, or when the generated code has not been thoroughly tested or verified. Even though the code is thoroughly reviewed, understood, and tested, self-declaration remains valuable for the ethical use of AI-generated code. Relevance. Our survey revealed that developers often self-declare AI-generated code mainly to support later debugging and review, driven by quality and security concerns. Even when code is modified or tested, many developers view self-declaration as an ethical practice that ensures transparency.
@@
Using simple self-declaration comments for better transparency. When adding AI-generated code to a project, developers may include a short self-declaration comment. This comment should note that the code was AI-generated, the name of the tool used, and the name of the code snippet (e.g., the name of the class, function, or language construct). Relevance. Our mining study showed that developers commonly include short self-declaration comments, and developers in the practitioners’ survey shared that self-declaring AI-generated code helps ensure transparency in development, promotes the ethical use of AI-generated code, and facilitates the distinction between AI-generated and human-written code.
@@
Adding code explanations for complex AI-generated code. A good practice for developers is to include a brief explanation alongside the simple self-declaration comments, especially when the AI-generated code is relatively complex. These comments should describe how the code works, why it is added, and any key details about the integrated AI-generated code that team members may need in the future. These code explanations should be kept concise, and developers should be mindful of the code’s readability. Relevance. Our mining study revealed that developers frequently include code explanations alongside self-declaration comments. Survey participants supported this practice, especially when the generated code is complex, as they believe it improves code maintainability and enhances knowledge sharing within the team.
@@
Include code quality indications in self-declaration comments. Developers should add details about the quality of AI-generated code in the self-declaration comments. This can include whether the code has been tested, whether it was modified by a human, and if there are any known issues or limitations in the AI-generated code. Including such information helps team members assess the code’s reliability and determine if further improvements are necessary. Relevance. In the mining study, we observed that developers often add comments about code quality, especially regarding whether AI-generated code has been tested or has known limitations. Additionally, survey participants also highlighted that quality concerns are a key reason for self-declaring AI-generated code.
@@
Document the context of AI-generated code. When using AI-generated code, developers should consider documenting its use context, such as the prompts used. Keeping a record of such information can help developers understand how the code was generated and make it easier to modify or troubleshoot in the future. Relevance. Our mining study identified several instances where developers included contextual information, such as prompts used during code generation.
@@
Using snippet-level comments for better traceability. It is advised to use snippet-level rather than file-level comments when self-declaring AI-generated code, as they provide a precise and more detailed approach to track AI-generated code snippets. However, if a file contains mostly AI-generated code or multiple AI-generated code snippets, file-level comments may be more suitable, as they offer a more concise summary without excessive repetition. Relevance. The mining study and survey responses revealed that developers predominantly use snippet-level comments for precise tracking of AI-generated code, but utilize file-level comments for efficiency of self-declaration when the file is mostly AI-generated.
@@ 6.2. Construct Validity
Keyword-based search. We used a set of search terms to identify and locate AI-generated code on GitHub repositories. This approach introduced a threat to validity as the search terms used might not be comprehensive or inclusive. To minimize this threat, we first conducted a pilot search using four search strategies and tested their corresponding groups of search terms (see Table 1), which helped us finalize the list of potential search terms. We iteratively added new terms to improve the coverage of our keyword. As a result, we identified a list of 23 search terms with broad coverage. The finalized 23 search terms include thanking terms, such as “Thanks ChatGPT”. These search terms introduced a threat to the validity because they could introduce bias in the study results. To minimize this threat in data collection, we also included warning search terms such as “warning Copilot” and also searched using these terms in addition to thanking terms (see Section 3.2.1). Secondly, we also ensured that all self-declaration comments appreciating AI tools are not included using thanking terms. Consequently, we have a large number of self-declaration comments appreciating AI tools, but they were not collected using thanking search terms.
@@ 7. Conclusions and Future Work
This study explores the practices of developers’ self-declaration of AI-generated code. We examined whether developers self-declare AI-generated code, their motivations for doing so, and the methods they use to achieve this. We employed a mixed-methods approach, combining data from GitHub repositories with a follow-up practitioners’ survey among practitioners. We found that developers have varying opinions on self-declaring AI-generated code, and their self-declaration practices are influenced by the proportion of AI contribution in the generated code. Most practitioners (76.6%) always or sometimes self-declare AI-generated code. In contrast, other practitioners (23.4%) noted that they never self-declare AI-generated code. Top reasons for self-declaring AI-generated code include (1) the need to track and monitor code for future review and debugging, and (2) ethical considerations related to the use of code generation tools. On the other hand, reasons for not self-declaring AI-generated code include (1) AI-generated code often requires extensive human modification, and (2) developers view self-declaration as an unnecessary activity, comparable to utilizing online documentation or coding forums. We then provided implications for researchers and a set of guidelines for developers to follow when self-declaring AI-generated code.
Our study also opens up several directions for future research. (1) We identify various reasons for self-declaring (or not) AI-generated code, which can be further investigated with a larger set of developers. Additionally, we provide guidelines for self-declaring AI-generated code by analyzing how developers report their generated code based on data from the mining study and practitioners’ survey. Future research can further refine and expand the guidelines using new datasets and validate their usefulness through empirical evaluation. (2) Our research shows that developers consider self-declaration comments beneficial for tracking AI-generated code for future review, debugging, and improvement. This opens up avenues for future research to investigate the impact of self-declaration comments on various code quality attributes, such as readability and maintainability. Such studies can contribute to a deeper understanding of the evolving landscape of software development, where AI-generated code is increasingly integrated into modern software systems. (3) It would be valuable to provide a code-commenting function in AI code generation tools to support the guidelines (see Section 5.3) for developers to declare AI-generated code.